Circular No. 8368

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2004cz IN ESO 407-G9 C. Jacques, Belo Horizonte, Brazil; and T. Napoleao, São Paulo, Brazil,

report the discovery of an apparent supernova (mag 15.0) on unfiltered CCD images taken with a 0.30-m reflector on June 26.33 and July 4.27 UT. The new object is located at α = 23^{h}12^{m}43^{s}.22, δ = −37^{o}12′58′′.7 (equinox 2000.0), which is 24′′.6 west and 30′′.4 south of the center of ESO 407-G9. Nothing is visible at this location on a CCD image taken on 2001 Oct. 17.09 (limiting mag 18.6) or on a red Palomar Survey image from 1980.

URANUS H. B. Hammel, Space Science Institute; and I. de Pater, University of

California at Berkeley, report the discovery that a discrete cloud feature in the southern hemisphere of Uranus has become detectable in K′-band adaptive-optics imaging on July 4.5 UT with the 10-m Keck II Telescope (+ NIRC2/AO) on Mauna Kea. This is the first known sighting of a southernhemispheric feature at this wavelength; all other 2.2-μm-bright Uranian features have been seen in the planet’s northern hemisphere. The feature was tracked for 3.5 hr as it transited the planetary disk, and throughout that time it remained co-located with the bright core of a large bright feature (lat. −38^{o} ± 2^{o}, long. ∼ 93^{o} ± 2^{o}) in J- and H-band images. Detection at K′ indicates that convective activity in the feature’s bright core has reached a higher altitude than has the activity in other southern features.

V5114 SAGITTARII D. K. Lynch, R. J. Rudy, S. Mazuk, and C. C. Venturini, Aerospace

Corporation; R. C. Puetter, University of California at San Diego; and R. B. Perry, Langley Research Center, NASA, report 0.4–2.5-μm spectroscopy of V5114 Sgr on June 22.4 UT using the enhanced Aerospace NIRIS spectrograph on the Lick Observatory 3-m telescope. The nova is well into its coronal phase, showing lines of [S VIII] 0.9911-μm, [Si IX] 1.2517-μm, and [Si VI] 1.9641-μm. [Ca VIII] 2.3205-μm was not seen. The Lyβ-fluoresced O I lines at 1.1287 and 0.8446 μm were strong. Also present were the unidentified novae lines at 1.1114, 1.1900, 1.5545, and 2.0996 μm. Balmer, Paschen, and Bracket lines are evident, and He I 1.0830-μm is the strongest line in the spectrum. The lines were flat-topped or slightly doubled, indicating spherical or nearly spherical expansion with FWHM ≈ 2200 km/s.

2004 July 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green